OpenStudy (anonymous):

What is the approximate circumference of a semicircle with a diameter of 64 meters? A. 100.48 m B. 200.96 m C. 401.92 m D. 1607.68 m

OpenStudy (anonymous):

@jzellem1004

OpenStudy (anonymous):

64*pi=200.96 so B is the correct answer

OpenStudy (ipwnbunnies):

No, it's a semicircle. A semicircle is half of a circle. So, the circumference will also be half of a full circumference.

OpenStudy (ipwnbunnies):

So, the diameter of the semicircle will actually be \[\frac{200.96 inches^2}{2}\]

OpenStudy (anonymous):

ah right so you'd take 200.96/2 = 100.48

OpenStudy (anonymous):

than you both your very helpful. but since I can only give 1 medal I will give one of you a medal and fan the other. if that is alright?